OpenStudy (anonymous):

What is the 5th term of an arithmetic sequence if t2 = -5 and t6 = 7?

OpenStudy (anonymous):

3 4 5 6

OpenStudy (anonymous):

t6-t2 = 4 * difference 7-(-5) = 4 * difference Difference = 3 t5 = t6 - difference = 7 - 3 = 4